A circular piece of cardboard with a diameter of 1 m will be made into a conical hat 40 cm high by cutting a sector off and joining the edges to form a cone. Determine the angle subtended by the sector removed.


1
Expert's answer
2021-05-30T22:16:01-0400


Given diameter of circular piece is 1 metre = 100 centimetres


So, radius = Diameter/2 = 100/2 = 50 centimetres


Given height of conical hat = 40 centimetres


From Pythagoras theorem,


r2 = h2 + x2


x2 = (50)2 - (40)2 = 2500-1600 = 900


x = 30 centimetres


Let C1 be circumference of circular piece


C2 be circumference of base of conical hat


C be length of arc


C = C1 - C2 = 2Ï€r - 2Ï€x = 2Ï€(50) - 2Ï€(30) = 40Ï€


C = r*θ


40π = 50*θ


θ = (40/50)π*(360°/2π) = 144°